# Topological superconductivity in quantum wires mediated by helical magnons

### v1

@@ Abstract
Triplet superconductivity arises in systems where the effective electronic attraction favors Cooper pairs with a symmetric spin structure, as may occur in non-collinear magnetic systems. Here we show that topological triplet superconductivity can be realized in a one-dimensional quantum wire through the interaction with magnetic fluctuations of a helical magnet. The magnon-mediated attraction favors triplet superconductivity over a large sector of magnetic parameters, and stabilizes a topological phase with a superconducting gap of the order of $1$ meV over an extended region of chemical potentials.
A promising path to realize fault-tolerant quantum computing is to utilize the excitations of topologically ordered systems Wen and Niu (1990); Kitaev (2003); Nayak et al. (2008). These excitations, generically named anyons Leinaas and Myrheim (1977); Wilczek (1982), are predicted to appear across a wide range of platforms, such as fractional quantum Hall systems, quantum spin liquids, and topological superconductors Arovas et al. (1984); Kitaev (2006); Savary and Balents (2016); Bartolomei et al. (2020); Nakamura et al. (2020). In particular, the sub-class of excitations called non-abelian anyons allow for information to be non-locally encoded and processed in the braiding patterns of the anyon world lines Kitaev (2003); Nayak et al. (2008). In topological superconductors, non-abelian anyons appear at the edges or in vortex cores of chiral two-dimensional superconductors Fu and Kane (2008); Nakosai et al. (2013); Sun et al. (2016); Ménard et al. (2017); Palacio-Morales et al. (2019); Wang et al. (2020); Kezilebieke et al. (2020), and at the ends of one-dimensional (1-D) quantum wires Flensberg (2010); Oreg et al. (2010); Lutchyn et al. (2010, 2018). However, due to the difficulties associated with realizing such phases, conclusive evidence of topological superconductivity has not been found in any experimental platform.
@@
It has recently been proposed that chiral topological superconductivity can be realized in two-dimensional heterostructures Yang et al. (2016); Rex et al. (2019); Bedow et al. (2020); Díaz et al. (2021); Mæland and Sudbø (2022, 2023); Mæland et al. (2023), where itinerant electrons interact with a non-collinear magnetic structure, or in chains of magnetic atoms deposited on a conventional superconductor Nadj-Perge et al. (2014); Neuhaus-Steinmetz et al. (2022); Schneider et al. (2023); Awoga et al. (2023). In 1-D systems, superconductors with triplet ($p$-wave) pairing are known to always reside in a topological class that can host non-trivial topological phases Altland and Zirnbauer (1997); Kitaev (2001); Chiu et al. (2016). To engineer topological superconductivity in 1-D, the dominant approach has been to combine proximity-induced $s$-wave superconductivity with strong spin-orbit coupling (SOC) semi-conductors and time-reversal symmetry breaking magnetic fields Fu and Kane (2008); Law et al. (2009); Stanescu et al. (2011); Alicea et al. (2011); Mourik et al. (2012); Das et al. (2012); Finck et al. (2013); Albrecht et al. (2016); Deng et al. (2016); Nichele et al. (2017); Zhang et al. (2019); Prada et al. (2020); Aghaee et al. (2023). Although straightforward in principle, this approach is in practice associated with a number of challenges, the most prominent being that the time-reversal breaking field must be finely tuned such that it is large enough to realize an effective spin-polarized regime, but small enough not to suppresses the effective superconducting gap Fu and Kane (2008); Oreg et al. (2010); Lutchyn et al. (2010). To avoid this problem it would be desirable to find a protocol for topological superconductivity where the spin-polarized regime and the superconducting gap both are increasing functions of some control parameter.
To address this challenge, we here propose a platform to stabilize triplet superconductivity by coupling a quantum wire to a helical magnet (see Fig. 1). The non-collinear magnetic order induces an effective SOC and Zeeman field among the electrons (see Fig. 2), thereby realizing an effective single-band regime. The magnon excitations of the helical magnet further mediate an effective attractive interaction between the electrons, which stabilizes triplet superconductivity with a superconducting gap of the order of $1$ meV over a large portion of the magnetic phase space. Within the effective single-band regime, the system enters a topological phase, with unpaired Majorana bound states at each end of the wire. Crucially, both the size of the effective single-band regime and the superconducting gap are increasing functions of the spin-electron coupling $g$. Our proposal thereby identifies quantum wires in proximity to helical magnets as a promising platform to explore topological superconductivity in 1-D.
@@
The competition between exchange and DMI stabilizes a helical magnetic order, characterized by a propagation vector ${\bf q}$ and a vector $\hat{\bf n}$ defining the plane of polarization. Specifically, the magnitude of the spiral momentum is given by $\sqrt{2}\tan q=D/J$, while the directions of the momentum and polarization vectors are determined by the normalized DMI vector $\hat{\bf D}$. Here we consider a DMI such that the spiral momentum ${\bf q}$ lies along the wire axis $\hat{\bf z}$, and with $\hat{\bf n}$ parallel to ${\bf q}$ (see Fig. 1). The equilibrium spin texture is then a spin helix, and can be written as ${\bf S}_{i}=\cos({\bf q}\cdot{\bf r}_{i}){\bf e}_{1}+\sin({\bf q}\cdot{\bf r}_{i}){\bf e}_{2}$, where the vectors ${\bf e}_{\alpha}$ and $\hat{\bf n}$ define a right-handed orthonormal system Sup .
@@
There are two main effects of the spin spiral on the electronic structure: First, the coupling to the equilibrium magnetic structure ${\bf S}_{i}$ induces an effective SOC and Zeeman splitting of the electronic bands (see Fig. 2a). Second, the coupling to magnon fluctuations around the helical configuration generates an effective attractive interaction among the electrons, which ultimately leads to superconductivity. The effect of the static spin spiral can be exactly accounted for, by diagonalizing the electronic sub-system in presence of the spiral ${\bf S}_{i}$ Viñas Boström et al. (2021); Sup , and results in the electron Hamiltonian $H_{e}=\sum_{k\tau}\epsilon_{k\tau}d_{k\tau}^{\dagger}d_{k\tau}$. Here $\tau$ denotes the electronic bands, and the momentum runs over the electronic Brillouin zone $[-\pi/a,\pi/a]$. The spiral shifts the minima of the originally spin-degenerate bands to ±$\pm q/2$ (Fig. 2a), thereby inducing an effective Rashba SOC. This shift is independent of the value of $g$, and is set for any finite $g$ by the momentum of the spin spiral. In addition, the coupling opens a gap of size $\Delta_{b}=2gS$ at $k=0$ and $k=\pm\pi/a$, acting like an effective Zeeman field. Together these effects realize an effective single-band regime for chemical potentials inside the gap.
@@
where $\beta=1/(k_{B}T)$ is the inverse temperature, and $\xi_{k}=\epsilon_{k}-\mu$. The gap equation can be viewed as an eigenvalue problem for the susceptibility matrix $\chi_{\tau_{1}\tau_{2}}^{\tau_{3}\tau_{4}}(k,k^{\prime})$, implicitly defined by the right hand side of Eq. 2, and superconductivity is signaled by the largest eigenvalue of this matrix exceeding unity.
@@
This qualitative analysis is in good agreement with the numerical solution of Eq. 2, showing that the dominant superconducting gap has triplet symmetry over a large fraction of the phase diagram (see Fig. 2b). Only in regions where $\epsilon_{u}(k_{Fu})-\epsilon_{d}(k_{Fd})\lesssim\Omega$, such as close to the ferromagnetic limit $q=0$, or when both $k_{F\tau}$ are close to the band edges (around $k_{F\tau}=0$), does the dominant gap have singlet symmetry. In the strict ferromagnetic limit, this result can be established analytically Sup , by noting that ferromagnetic magnons only scatter electrons with opposite spins. That such a large portion of the phase diagram is dominated by triplet pairing follows from the fact that singlet pairing requires inter-band scattering. For $k_{Fd}\neq k_{Fu}$, such scattering results in Cooper pairs with a finite center-of-mass momentum $k_{Fu}-k_{Fd}\sim q$, which strongly suppresses singlet pairing at finite $q$ Loder et al. (2010). With increasing $q$, intra-band magnon scattering therefore quickly becomes the dominant pairing mechanism, resulting in a gap with triplet symmetry. We note that in the antiferromagnetic limit $q=\pi$, an effective gap opens between the electronic bands, and only triplet pairing is possible.
@@
[…]
We have demonstrated that magnon fluctuations of a helical magnet can mediate triplet superconductivity in a proximate quantum wire. As the induced band gap and superconducting gap both increase with the spin-electron coupling $g$, various material engineering techniques such as wire geometry optimization, microstructuring or cavity-enhanced light-matter couplings can be employed to maximize its value Moll (2018); Sentef et al. (2018); Nilsson et al. (2018); Barrigón et al. (2019); Viñas Boström et al. (2023). We expect that, utilizing such strategies, an effective coupling exceeding $g\approx 50$ meV can be achieved, allowing our protocol to surpass present state-of-the-art experiments with a topological superconducting gap of the order $\Delta\sim 0.1$ meV Aghaee et al. (2023). The present model is a simplified description of a real quantum wire, and in reality effects coming from repulsive electron-electron interactions, disorder and a more complex band structure should be included Das Sarma (2023). While the former is expected to compete with the magnon-mediated attraction, the later could result in a larger DOS at the Fermi level, thereby enhancing superconductivity. In conclusion, our results identify quantum wires in proximity to non-collinear magnets as a promising platform to explore topological superconductivity in 1-D.

### v2

# Magnon-mediated topological superconductivity in a quantum wire

@@ Abstract
Many emergent phases of matter stem from the intertwined dynamics of quasi-particles. Here we show that a topological superconducing phase emerges as the result of interactions between electrons and magnons in a quantum wire and a helical magnet. The magnon-mediated interaction favors triplet superconductivity over a large magnetic phase space region, and stabilizes topological superconductivity over an extended region of chemical potentials. The superconducting gap depends exponentially on the spin-electron coupling, allowing it to be enhanced through material engineering techniques.
The intertwined dynamics of collective excitations underlies many diverse phenomena observed in condensed matter systems. A prominent example where such interactions play a key role is in the effective phonon-mediated attraction between electrons, responsible for the superconducting instability of simple metals at low temperatures Bardeen et al. (1957). Other emergent phases driven by quasi-particle interactions include the spontaneous electric polarization of ferroelectric and multiferroic materials in response to lattice distortions Eerenstein et al. (2006); Li et al. (2019), the binding between electrons and holes and their possible subsequent condensation into an excitonic insulator Werdehausen et al. (2018); Mazza et al. (2020), and the stabilization of fractional quantum Hall states through electron-electron interactions Laughlin (1983).
Many emergent phases themselves give rise to quasi-particles with potentially unconventional properties. In particular, the excitations of topologically ordered systems Wen and Niu (1990); Kitaev (2003); Nayak et al. (2008) such as fractional quantum Hall systems, quantum spin liquids, and topological superconductors Arovas et al. (1984); Kitaev (2006); Savary and Balents (2016); Bartolomei et al. (2020); Nakamura et al. (2020), termed anyons Leinaas and Myrheim (1977); Wilczek (1982), have garnered much interest due to their promise in realizing fault-tolerant quantum computing. In particular, the sub-class known as non-abelian anyons allow for information to be non-locally encoded and processed in the braiding patterns of the anyon world lines Kitaev (2003); Nayak et al. (2008). In topological superconductors, non-abelian anyons may appear in vortex cores of chiral two-dimensional superconductors Fu and Kane (2008); Nakosai et al. (2013); Sun et al. (2016); Ménard et al. (2017); Palacio-Morales et al. (2019); Wang et al. (2020); Kezilebieke et al. (2020), or at the ends of one-dimensional (1-D) quantum wires Flensberg (2010); Oreg et al. (2010); Lutchyn et al. (2010, 2018); Law et al. (2009); Stanescu et al. (2011); Alicea et al. (2011); Mourik et al. (2012); Das et al. (2012); Finck et al. (2013); Albrecht et al. (2016); Deng et al. (2016); Nichele et al. (2017); Zhang et al. (2019); Prada et al. (2020); Aghaee et al. (2023). However, to stabilize a topological superconducting phase, the pairing needs to be mediated by quasi-particles carrying an intrinsic spin structure that favors spin triplet over spin singlet pairing.
@@
As a prominent example, magnon-mediated topological superconductivity has recently been proposed for two-dimensional systems with non-collinear magnetic order, such as skyrmion crystals and helical magnets in proximity to a two-dimensional normal metal surface Mæland and Sudbø (2022, 2023); Mæland et al. (2023). However, since the majority of experimental work on topological superconductivity is concerned with one-dimensional structures, identifying criteria for realizing magnon-mediated topological superconductivity in one dimension is of key importance. Indeed, both the superconducting and topological phenomenology is qualitatively different in one and two dimensions, stemming in large part from the different symmetry requirements necessary to realize a non-trivial topology in different dimensions Altland and Zirnbauer (1997); Kitaev (2001); Chiu et al. (2016). In particular, to stabilize topological superconductivity in 2-D it is necessary to obtain a time-reversal symmetric and fully gapped superconducting state, while in 1-D systems a topological superconducting phase is realized by breaking time-reversal symmetry (thereby realizing an effective single-band regime), while simultaneously stabilizing triplet pairing. As will be demonstrated below, both these conditions are satisfied by the dynamical coupling between itinerant electrons and the magnons of a helical magnet, away from the collinear ferromagnetic and antiferromagnetic limits.
Specifically, we here investigate superconductivity resulting from the coupling between a quantum wire and a helical magnet (see Fig. 1). The magnons of the helical magnet mediate an effective attraction between electrons of arbitrary spin projection, thereby stabilizing unconventional triplet superconductivity over a large region of phase space. The non-collinear magnetic order induces an effective spin-orbit coupling (SOC) and Zeeman field among the electrons, that allows to realize an effective single-band regime over a finite range of chemical potentials. Within the single-band regime, the system enters a topological phase, with unpaired Majorana bound states at each end of the wire. Crucially, both the size of the effective single-band regime and the superconducting gap are increasing functions of the spin-electron coupling $g$. Our proposal thereby identifies quantum wires in proximity to helical magnets as a promising platform to realize topological superconductivity, without the need to proximitize the wire to a conventional superconductor. It also showcases the power of utilizing quasi-particle interactions to realize unconventional forms of matter.
@@
The competition between exchange and DMI stabilizes a helical magnetic order, characterized by a propagation vector ${\bf q}$ and a vector $\hat{\bf n}$ defining the plane of polarization. Specifically, the magnitude of the spiral momentum is given by $\tan q=D/J$, while the directions of the momentum and polarization vectors are determined by the normalized DMI vector $\hat{\bf D}$. Here we consider a DMI such that the spiral momentum ${\bf q}$ lies along the wire axis $\hat{\bf z}$, and with $\hat{\bf n}$ parallel to ${\bf q}$ (see Fig. 1). The equilibrium spin texture is then a spin helix, and can be written as ${\bf S}_{i}=\cos({\bf q}\cdot{\bf r}_{i}){\bf e}_{1}+\sin({\bf q}\cdot{\bf r}_{i}){\bf e}_{2}$, where the vectors ${\bf e}_{\alpha}$ and $\hat{\bf n}$ define a right-handed orthonormal system Sup .
@@
There are two main effects of the spin spiral on the electronic structure: First, the coupling to the equilibrium magnetic structure ${\bf S}_{i}$ induces an effective SOC and Zeeman splitting of the electronic bands (see Fig. 2a). Second, the coupling to magnon fluctuations around the helical configuration generates an effective attractive interaction among the electrons, which ultimately leads to superconductivity. The effect of the static spin spiral can be exactly accounted for, by diagonalizing the electronic sub-system in presence of the spiral ${\bf S}_{i}$ Viñas Boström et al. (2021); Sup , and results in the electron Hamiltonian $H_{e}=\sum_{k\tau}\epsilon_{k\tau}d_{k\tau}^{\dagger}d_{k\tau}$. Here $\tau$ denotes the electronic bands, and the momentum runs over the electronic Brillouin zone $[-\pi/a,\pi/a]$. The spiral shifts the minima of the originally spin-degenerate bands to ±$\pm q/2$ (Fig. 2a), thereby inducing an effective Rashba SOC. This shift is independent of the value of $g$, and is set for any finite $g$ by the momentum of the spin spiral. In addition, the coupling opens a gap of size $\Delta_{b}=2gS$ at $k=0$ and $k=\pm\pi/a$, acting like an effective Zeeman field. Together these effects realize an effective single-band regime for chemical potentials inside the gap. The static effects of our model have been considered within the context of so-called Yu-Shiba-Rusinov chains Nadj-Perge et al. (2013), which give rise to a similar induced spin-orbit coupling and Zeeman splitting of the electronic bands. However, these earlier treatments neglect the dynamical effects of the spin-electron interaction, which here gives rise to an intrinsic triplet superconductivity.
@@
where $\beta=1/(k_{B}T)$ is the inverse temperature, and $\xi_{k}=\epsilon_{k}-\mu$. The gap equation can be viewed as an eigenvalue problem for the susceptibility matrix $\chi_{\tau_{1}\tau_{2}}^{\tau_{3}\tau_{4}}(k,k^{\prime})$, implicitly defined by the right hand side of Eq. 2, and superconductivity is signaled by the largest eigenvalue of this matrix exceeding unity Allen and Mitrović (1983).
@@
This qualitative analysis is in good agreement with the numerical solution of Eq. 2, showing that the dominant superconducting gap has triplet symmetry over a large fraction of the phase diagram (see Fig. 2c). Only in regions where $\epsilon_{u}(k_{Fu})-\epsilon_{d}(k_{Fd})\lesssim\Omega$, such as close to the ferromagnetic limit $q=0$, or when both $k_{F\tau}$ are close to the band edges (around $k_{F\tau}=0$), does the dominant gap have singlet symmetry. In the strict ferromagnetic limit, this result can be established analytically Sup , by noting that ferromagnetic magnons only scatter electrons with opposite spins. That such a large portion of the phase diagram is dominated by triplet pairing follows from the fact that singlet pairing requires inter-band scattering. For $k_{Fd}\neq k_{Fu}$, such scattering results in Cooper pairs with a finite center-of-mass momentum $k_{Fu}-k_{Fd}\sim q$, which strongly suppresses singlet pairing at finite $q$ Loder et al. (2010). With increasing $q$, intra-band magnon scattering therefore quickly becomes the dominant pairing mechanism, resulting in a gap with triplet symmetry. We note that in the antiferromagnetic limit $q=\pi$, an effective gap opens between the electronic bands, and only triplet pairing is possible.
@@
To realize the model of Eq. 1, several strategies are possible. The most straightforward is to consider a quantum wire in proximity to a helical magnet, as illustrated in Fig. 1. In this case, the wire could consist either of a simple metal, or a lightly doped or gate tuned semi-conductor (such as InAs Nilsson et al. (2018); Debbarma et al. (2022)), with a single active and largely uncorrelated band. For the helical magnet, it is preferable to use a material with a short spiral wave length ($\lambda\sim 10$ nm), low magnon energy and large spin length, such as MnGe, MnSi or NiI2 Nagaosa and Tokura (2013); Song et al. (2022). To maximize $g$, it is desirable to use a wire geometry where as much as possible of the wire is strongly coupled to the magnet. Estimating $g$ for a general interface is hard, but based on earlier work it is expected that $g\sim 10-100$ meV can be achieved Qiao et al. (2014); Norden et al. (2019); Mæland et al. (2021); Cardoso et al. (2023). An alternative strategy is to consider a quasi-one-dimensional multi-band system Schlappa et al. (2012); Kennes et al. (2020), where the magnetic order and itinerant electrons co-exist within the same material. In this case the spin-electron coupling is effectively the Hund’s coupling $J$, potentially leading to large effective $g$ on the order of $1$ eV. However, since such multi-band systems are typically strongly correlated, Eq. 1 should in this case be supplemented with additional interaction terms.
We have demonstrated that magnon fluctuations of a helical magnet can mediate triplet superconductivity in a proximate quantum wire. As the induced band gap and superconducting gap both increase with the spin-electron coupling $g$, various material engineering techniques such as wire geometry optimization, microstructuring or cavity-enhanced light-matter couplings can be employed to maximize its value Moll (2018); Sentef et al. (2018); Nilsson et al. (2018); Barrigón et al. (2019); Viñas Boström et al. (2023). We expect that, utilizing such strategies, an effective coupling exceeding $g\approx 50$ meV can be achieved, allowing our protocol to surpass present state-of-the-art experiments with a topological superconducting gap of the order $\Delta\sim 0.1$ meV Aghaee et al. (2023). Compared to the conventional scheme of generating 1-D topological superconductivity combining proximity-induced superconductivity with large Rashba spin-orbit coupling and external magnetic fields, the main advantage of our proposal is the absence of fine tuning conditions.
The present model is a simplified description of a real quantum wire, and in reality effects coming from repulsive electron-electron interactions, disorder and a more complex band structure should be included Das Sarma (2023). While the former is expected to compete with the magnon-mediated attraction, the later could result in a larger DOS at the Fermi level, thereby enhancing superconductivity. Another route to connect the present model with more realistic experimental setups is through the study of the corresponding system with open boundary conditions. In conclusion, our results identify quantum wires in proximity to non-collinear magnets as a promising platform to explore topological superconductivity in 1-D.